EDITORIAL

Terry Goodman

I noted with interest the range of topics and discussions represented in the program from the MAA–AMS Joint Meeting held in New Orleans this past January. In particular, I was struck by the number of sessions/discussions focusing on mathematics education reform. Most notable were the special sessions co-sponsored by AMS, MAA, and MER. The Mathematicians and Education Reform (MER) Forum is an organization that has consistently sought to encourage the mathematics community to engage in a variety of activities that will promote thoughtful discussion and recognize exemplary programs and projects.

To quote from some of MER’s materials, “The Mathematicians and Education Reform (MER) Forum seeks the effective participation of mathematicians in mathematics education reform at the K-12, undergraduate, and graduate levels, and the recognition of the importance of these efforts to the well being of the mathematics community.”

MER encourages active participation by providing a variety of workshops and programs that it offers throughout the year. As mentioned previously, at this year’s joint meetings MER coordinated a special session of Mathematics and Education Reform. Further, MER publishes a variety of materials, including the MER Newsletter, MER eNews, and several volumes of collected articles.

Addressing the many problems and issues associated with mathematics education reform will require broad participation from a vareity of groups. Mathematicians can provide valuable input relative to curricular reform, pedagogical issues, and teacher preparation and development. MER provides a forum that allows the mathematics community to engage in important discussions and develop relevant activities and materials.

If you would like more information concerning MER, please consult its Home Page at \underline{http://www.math.uic.edu/MER}.